OpenStudy (anonymous):

Which is an equivalent form of 12/15? 1 1/4 24/30 4/7 36/39

OpenStudy (dangerousjesse):

What's \(\large\frac{12}{15}\times\frac{2}{2}\)?

OpenStudy (anonymous):

4/5

OpenStudy (dangerousjesse):

Don't simplify it :)

OpenStudy (anonymous):

Huh?

OpenStudy (dangerousjesse):

I meant for you to multiply the numerator and denominator by two, and leave it be.

OpenStudy (anonymous):

What would be the answer

OpenStudy (dangerousjesse):

You tell me

OpenStudy (anonymous):

C:-)

OpenStudy (dangerousjesse):

No, not even close :P Just cross multiply 12/15 by two to get the answer.

OpenStudy (anonymous):

I got 8/5

OpenStudy (dangerousjesse):